# The Gray graph is pseudo 2-factor isomorphic

### v1

@@ Abstract
Keywords. Cubic, bipartite, 2-factor, conjecture, counterexample, computation, Gray graph
@@ 1 Introduction
[18, Conjecture $3.2$] Let $G$ be a $2$-factor Hamiltonian $k$-regular bipartite graph. Then either $k=2$ and $G$ is a cycle or $k=3$ and $G$ can be obtained from $K_{3,3}$ and $H_{0}$ by repeated star products.
@@
In a cubic graph, the three edges incident with a vertex constitute a $3$-edge-cut because their removal leaves an isolated vertex, and is called trivial, other edge-cuts being non-trivial. A cubic graph is said to be essentially $4$-edge-connected if it does not have non-trivial $3$-edge-cuts. The star product of graphs has an inverse operation called 3-cut reduction which consists of removing a non-trivial $3$-edge-cut and adding a vertex of degree 3 to each of the two remaining components. By repeatedly performing this operation on a cubic graph $G$, we arrive at the essentially $4$-edge-connected constituents of $G$.
A set $S$ of edges of a graph $G$ is said to be a cyclic edge-cut if removing $S$ from $G$ results in two components, each containing a cycle. A graph $G$ is said to be cyclically $m$-edge-connected if every cyclic edge-cut in $G$ has a size of at least $m$. For an integer $k$ with $k\leq 4$, a cubic graph is essentially $k$-edge-connected if and only if it is cyclically $k$-edge-connected (see e.g. [11]). Please note that graphs obtained from $K_{3,3}$ and $H_{0}$ by repeated star products are not cyclically 4-edge-connected because of the $3$-edge-cut produced by the definition of the star product.
@@
Recently in [20] it was proved that Conjecture 1.1 is also equivalent to [20, Conjecture 1.8]: The Heawood graph and $K_{3,3}$ are the only $2$-factor Hamiltonian, cubic braces. Moreover, it can be narrowed to [20, Conjecture 1.9]: All non-Pfaffian, cubic braces of girth at least $6$, of order congruent to $2$ modulo $4$, are not $2$-factor Hamiltonian. For definitions of braces and Pfaffian, please refer to [20].
Another equivalence of Conjecture 1.1 appears in the context of extending perfect matchings to Hamiltonian cycles and is stated in [34, Conjecture 7] as: Every bipartite cyclically 4-edge-connected cubic Perfect Matching Hamiltonian graph with girth at least 6, except the Heawood graph, admits a perfect matching which can be extended to a Hamiltonian cycle in exactly one way. For definitions of perfect matching and Perfect Matching Hamiltonian graph please refer to [34].
@@
The type of a $2$-factor in a graph is the tuple listing the lengths of its cycles in increasing order. So, for example the unique $2$-factor types of $K_{4}$, $K_{5}$, $K_{3,3}$, $H_{0}$ and the Petersen graph are respectively $(4),\,(5),\,(6),\,(14)$ and $(5,5)$; while the Pappus graph has $2$-factors of two types, precisely $(18)$ and $(6,6,6)$.
@@
Another possible drawing of $\mathcal{G}$ is shown in Figure 4, where the specific join between the two aforementioned configurations is highlighted and it also allows to explain its automorphism group.
@@ 3.2 Investigation of symmetrical graph censuses
Using a computer program (see Section 3.3 for details), we tested whether these censuses of highly symmetrical cubic graphs contain pseudo 2-factor isomorphic graphs different from the known examples. We did not find any new pseudo 2-factor isomorphic, nor any new 2-factor Hamiltonian graphs as outlined in 3.5 and 3.6.

### v2

@@ Abstract
Keywords. Cubic graph, bipartite graph, 2-factor, computation, Gray graph
@@ 1 Introduction
[17, Conjecture $3.2$] Let $G$ be a $2$-factor Hamiltonian $k$-regular bipartite graph. Then either $k=2$ and $G$ is a cycle or $k=3$ and $G$ can be obtained from $K_{3,3}$ and $H_{0}$ by repeated star products.
@@
In a cubic graph, the three edges incident with a vertex constitute a $3$-edge-cut because their removal leaves an isolated vertex, and is called trivial, other edge-cuts being non-trivial. A set $S$ of edges of a graph $G$ is said to be a cyclic edge-cut if removing $S$ from $G$ results in two components, each containing a cycle. A graph $G$ is said to be cyclically $m$-edge-connected if every cyclic edge-cut in $G$ has size at least $m$. A cubic graph is said to be essentially $4$-edge-connected if it is $3$-connected and every $3$-edge-cut is trivial. For an integer $k$ with $k\leq 4$, a cubic graph is essentially $k$-edge-connected if and only if it is cyclically $k$-edge-connected (see e.g. [10]). Please note that graphs obtained from $K_{3,3}$ and $H_{0}$ by repeated star products are not cyclically 4-edge-connected because of the $3$-edge-cut produced by the definition of the star product.
The star product of graphs has an inverse operation called 3-cut reduction which consists of removing a non-trivial $3$-edge-cut and adding a vertex of degree 3 to each of the two remaining components. By repeatedly performing this operation on a cubic graph $G$, we arrive at the essentially $4$-edge-connected constituents of $G$.
@@
Recently in [19] it was proved that Conjecture 1.1 is also equivalent to [19, Conjecture 1.8]: The Heawood graph and $K_{3,3}$ are the only $2$-factor Hamiltonian, cubic braces. Moreover, it can be narrowed to [19, Conjecture 1.9]: All non-Pfaffian, cubic braces of girth at least $6$, of order congruent to $2$ modulo $4$, are not $2$-factor Hamiltonian. For definitions of braces and Pfaffian, please refer to [19].
Another equivalence of Conjecture 1.1 appears in the context of extending perfect matchings to Hamiltonian cycles and is stated in [32, Conjecture 7] as: Every bipartite cyclically 4-edge-connected cubic Perfect Matching Hamiltonian graph with girth at least 6, except the Heawood graph, admits a perfect matching which can be extended to a Hamiltonian cycle in exactly one way. For definitions of perfect matching and Perfect Matching Hamiltonian graph please refer to [32].
@@
The type of a $2$-factor in a graph is the tuple listing the lengths of its cycles in increasing order. So, for example the unique $2$-factor types of $K_{4}$, $K_{5}$, $K_{3,3}$, $H_{0}$ and the Petersen graph are respectively (4),(5),(6),(14)(4),\,(5),\,(6),\,(14) and $(5,5)$; while the Pappus graph has $2$-factors of two types, precisely $(18)$ and $(6,6,6)$.
@@
Another possible drawing of $\mathcal{G}$ is shown in Figure 4, where the specific join between the two aforementioned configurations is highlighted and it also allows to explain its automorphism group. More alternative drawings are shown on Wolfram MathWorld [33].
@@ 3.2 Investigation of symmetrical graph censuses
Using a computer program (see Section 3.3 for details), we tested whether these censuses of highly symmetrical cubic graphs contain pseudo 2-factor isomorphic graphs different from the known examples. We did not find any new pseudo 2-factor isomorphic, nor any new 2-factor Hamiltonian graphs as outlined in ˜3.5 and 3.6. Recall that a graph $G$ is a Cayley graph if there exists a group $\Gamma$ and an inverse-closed subset $S\subseteq\Gamma\setminus\{1_{\Gamma}\}$ such that the vertices of $G$ are the elements of $\Gamma$, and two vertices $g,h\in\Gamma$ are adjacent whenever $h=gs$ for some $s\in S$. In this case we denote $G=Cay(\Gamma,S)$.